A mother wants to invest $6,000.00 for her son's future education. She invests a portion of the money in a bank certificate of deposit which earns 4% interest and the remainder in a savings bond that earns 7%. If the total interest earned after one year is $360.00, how much money was invested into the Certificate of Deposit account?
Answer by mananth(16946)

Investment I 4.00% per annum
Investment II 7.00% per annum

x+y 6000 ------------------------1
4.00%x+7.00% y=$360.00
Multiply by 100
4x+7y=$36,000.00 --------2

Multiply (1) by -4
we get

-4x-4y= -24000.00

Add this to (2)

3y = $12,000.00

divide by 3

y = $4,000.00 investment at 7.00%
Balance $2,000.00 investment at 4.00%
CHECK
$2,000.00 $80.00
$4,000.00 $280.00
Total $360.00